:49 It says: Solve by completing the square
equation: m^2-25/2m=5/2
the 25/2 is a fraction and the "x" is on the side, and the 5/2 is a fraction as well.
This is what I've tried so far:
I moved the "5/2" to the other side of the equal sign to where it looks like this:
m^2-25/2x-5/2=0
then I've tried to figure out how to square it

Answer by scott8148(6628)

m^2 - (25/2)m = 5/2

adding half of the 1st order coefficient, squared ___ m^2 - (25/2)m + (-25/4)^2 = 5/2 + (-25/4)^2

[m - (25/4)]^2 = 665/16

taking square root ___ m - (25/4) = ±[sqrt(665)]/4

m = [25 ± sqrt(665)]/4